# Bird’s-Eye View to Street-View: A Survey

### v1

@@ II-G1 CrossMLP [9]
TABLE II: Summary of the datasets used for Satellite to Street view translation.
@@
Dayton 2016 - - 354$\times$354 354$\times$354 55,000 21,048 76,048
CVUSA 2017 360 No 1232$\times$224 750$\times$750 35,532 8,884 44,416
CVACT 2019 360 Yes 1664$\times$832 1200$\times$1200 35,532 92,802 128,334
III Datasets
We discussed the primary three datasets used for satellite-to-street translation as given in Table II.
Dayton [19]: is introduced for ground-to-aerial or aerial-to-ground image translation or cross-view image synthesis, consisting of 76,048 pairs of satellite and street-view images across 11 different cities in the USA, where 55,000 are reserved for training and the rest for testing. Street-view images are collected from Google Maps of the USA, then several crops are made for each street-view panorama, and for each crop, a satellite image is queried using Google Maps with a resolution of 354$\times$354. Figure 9 showcases sample images from the dataset.
@@ III Datasets
CVUSA [2]: is an abbreviation of the Crossview United States of America for cross-view geo-localization tasks. The dataset comprises 44,416 satellite and panorama street-view images across the USA, with 35,532 pairs for training and 8,884 for testing. The images are collected from Flicker and Google Street View, having a resolution of 1232 $\times$ 224. For Flicker, the area has been divided into 100$\times$100 grid; 150 images from each grid cell have been downloaded from 2012 onwards, while the pictures for Google Street View are sampled from randomly selected places within the USA. Moreover, the satellite images are obtained from Bing Maps with a resolution of 750 $\times$ 750, where for each street-view image, an 800$\times$800 satellite image is downloaded centered on the exact location. Figure 10 showcases sample images from the dataset.
@@
CVACT [18]: stands for Australian Capital Territory City-Scale Cross-View, consisting of 128,334 pairs of satellite and panorama street-view images across Canberra City in Australia. The dataset has 35,532 for training and 92,802 for testing. Street-view images have been collected from the Google Street View API with an image resolution of 1664$\times$832. The authors have obtained the satellite images from the Google Map API. Further, the matched satellite image at the GPS position has been downloaded for each street-view image with a resolution of 1200$\times$1200. Figure 11 showcases sample images from the dataset.
@@ IV Evaluation Metrics
    Table I shows a few comparative results (i.e., accuracy and inception scores) using Conditional GANs on the CVUSA dataset.
@@
V challenges, limitations and Future directions
@@
The following sections highlight the challenges and limitations in the Satellite to Street View generation with possible future directions.
@@
Lack of multi-modalities: Using a single data modality depending on a single image, like a satellite image, for generating street view is more challenging than using multiple data modalities. However, some of the recent works [12, 13] rely on some auxiliary information like segmentation maps for generating street-view. Regardless, generating segmentation maps may limit the methods if there is any difficulty with its computation, affecting the quality of the synthesized street-view image. Multi-modal datasets might solve the problem and increase the quality of the generation.
@@
Lack of novel techniques: The methods used in the literature need to be updated; currently, generative adversarial network and their variants are the most popular choices for a street-view generation. Moreover, various CNNs are employed for feature extraction, while Siamese and Triplet Networks have been used for street-view geo-localization tasks. Researchers must consider and focus on recent state-of-the-art techniques such as transformers and stable diffusion for satellite to street-view synthesizing.
@@
VI Conclusion
This paper has surveyed the satellite to street-view synthesis domain from different aspects, including the recently available methods, common datasets (benchmark datasets), common evaluation metrics, and challenges and limitations, by examining 20 recently published papers. It is clear from reviewed papers that the Generative Neural Network (GAN) and its varieties are very immersed and widely used in this domain, despite other novel methods are available. Also, despite the success of the review methods, the methods in the literature still failed to generate small and detailed objects, which confirms that novel and updated methods need to be used for synthesizing more realistic and detailed street-view images. Along with this, a limited number of datasets have been used in the majority of the surveyed papers in this domain, which indicates the need for more publicly available datasets to enrich the research in this domain. In addition, more specific evaluation metrics need to be investigated and applied in this domain, as the methods in the literature have been evaluated on image quality-based metrics. Overall, the disclosed research challenges highlight that the satellite to street-view synthesis is still an active area of research.

### v2

@@ 3 Datasets
We discussed the primary three datasets used for satellite-to-street translation as given in Table 2.
Table 2: Summary of the datasets used for Satellite to Street view translation.
@@
Dayton 2016 - - $\times354$ $\times354$ 55,000 21,048 76,048
[…]
Satellite Image Refer to caption Refer to caption
Ground truth Refer to caption Refer to caption
Pix2pix [18] Refer to caption Refer to caption
X-Fork [36] Refer to caption Refer to caption
X-Seq [36] Refer to caption Refer to caption
SelectionGAN [46] Refer to caption Refer to caption
GeoUrban [32] Refer to caption Refer to caption
PAGAN [54] Refer to caption Refer to caption
Figure 12: Qualitative comparison on CVUSA dataset.
Satellite Image Refer to caption Refer to caption
Ground truth Refer to caption Refer to caption
Pix2pix[18] Refer to caption Refer to caption
X-Frok [36] Refer to caption Refer to caption
X-Seq [36] Refer to caption Refer to caption
SelectionGAN [46] Refer to caption Refer to caption
GeoUrban [32] Refer to caption Refer to caption
PAGAN [54] Refer to caption Refer to caption
Figure 13: Qualitative comparison on OP dataset.
Satellite Image Refer to caption Refer to caption
Ground truth Refer to caption Refer to caption
Pix2pix [18] Refer to caption Refer to caption
X-Frok [36] Refer to caption Refer to caption
SSVG [42] Refer to caption Refer to caption
Figure 14: Qualitative comparison on CVACT dataset.
Satellite Image ControlNet [58] CrossViewDiff [26] Ground truth
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 15: Qualitative comparison on CVACT dataset.
Dayton dataset [50] is introduced for ground-to-aerial or aerial-to-ground image translation or cross-view image synthesis, consisting of 76,048 pairs of satellite and street-view images across 11 different cities in the USA, where 55,000 are reserved for training and the rest for testing. Street-view images are collected from Google Maps of the USA, then several crops are made for each street-view panorama, and for each crop, a satellite image is queried using Google Maps with a resolution of $\times354$. Figure 16 showcases sample images from the dataset.
@@
CVUSA [53] dataset is an abbreviation of the Crossview United States of America for cross-view geo-localization tasks. The dataset comprises 44,416 satellite and panorama street-view images across the USA, with 35,532 pairs for training and 8,884 for testing. The images are collected from Flicker and Google Street View, having a resolution of 1232 $\times$ 224. For Flicker, the area has been divided into $\times100$ grid; 150 images from each grid cell have been downloaded from 2012 onwards, while the pictures for Google Street View are sampled from randomly selected places within the USA. Moreover, the satellite images are obtained from Bing Maps with a resolution of 750 $\times$ 750, where for each street-view image, an $\times800$ satellite image is downloaded centered on the exact location. Figure 17 showcases sample images from the dataset.
@@
[…]
@@ 4 Evaluation Metrics
    Table 3 shows a few comparative results (i.e., accuracy and inception scores) using Conditional GANs on the CVUSA dataset.
@@
5 Qualitative Comparison
Figure 12 (a) shows a qualitative comparison on the CVUSA dataset. It can be observed that Pix2pix and X-Seq trying to generate images that closely mimic the outlines of the ground truth image. In contrast, X-Fork and GeoUrban generate images with more artifacts compared to Pix2pix and X-Seq. Also, SelectionGAN demonstrates improvements in generating more clearly shaped streets and street lines, while the PAGAN method generates the most realistic street-view images compered to others. On the other hand, Figure 12(b) shows another qualitative example, clearly showing that X-Fork method generates low-quality images with noticeable artifacts. The Pix2pix and X-Seq methods generate better street-view images compared to X-Fork; however, artifacts are still present, and the details poorly defined. Whereas, SelectionGAN generates better images, by incorporating finer street lines and generating better more realistic trees appearance. PAGAN produces superior images in terms of quality and details. It incorporates more details appearance of trees, and improved street lines, bringing its generated image closer to the ground truth compared to other methods. In Figure 13 we provide a comparison on the OP dataset. As shown in Figure 13 (a), the Pix2pix, X-Fork methods suffer from a high-level of artifacts and generate noisy images. While X-Seq and SelectionGAN also exhibit noticeable artifacts, they manage to capture part of building on the right side of the image. The GeoUrban method generates a better generated image in terms of image outlines, making it slightly better-performing than other methods in this regard. PAGAN produces the best generated image in terms of outlines and boundaries, showing improved spatial consistency with the ground truth. However, in the central pert of the generated image, it incorrectly generate trees where a building appear in the ground truth. However, in overall, it generates more realistic street-view image in comparison to the other methods. In Figure 13 (b), the Pix2pix, X-Fork, and SelectionGAN methods generating artifact-heavy and noisy street-view images. In contrast, the X-Seq and GeoUrban methods generate more defined outlines and capture some street lines. PAGAN outperforms other methods by generating geometrically consistent street-view images, despite that some local details are missing in the output, it successfully include finer street lines compared to the other methods. Figures 14 show a qualitative comparison on the CVACT dataset. As shown in Figure 14(a), X-Forks generates a street-view image with noticeable artifacts and poorly shaped objects’ outlines. While Pix2pix generates better image compared to X-Fork; however it lacks full geometric consistency. SSVG method generates geometrically consistent street-view images with finer details compared with other methods. In Figure 14(b) the Pix2pix and X-Fork methods generate images with an acceptable level of geometric consistency and local details, such as street lines. In addition SSVG is able to generate the sidewalk in the middle, bringing it closer to the ground truth than other methods. On the other hand, Figure 15 (a) shows that ControlNet adding details that do not exists in the ground truth, making it deviate from the ground truth. In comparison, the CrossViewDiff method generates a geometrically consistent street-view image. Figure 15(b) shows that while CrossViewDiff generates a street-view image with different sidewalk shape on the right side of the image, it remains almost geometrically consistent, compared to the image generated by ControlNet, which represents a different street layout and introduces details not exist in the ground truth. In Figure 15 (c) Controlnet generates buildings, cars, and trees in locations where they do not exist in the ground truth, while CrossViewDiff is able to generate geometrically consistent street-view with finer-details. From this compression, we can conclude that the diffusion-based methods are capable to generate street-view images with a good level of geometrically consistency and fine-grained details. However, they may sometimes over-detail the output images, as seen with ControlNet. While ControlNet is a general-purpose diffusion model and not discussed in our methods section, we present its qualitative results here for the sake of completeness. It was reported in [26] within the context of satellite-to-street-view generation.
6 Quantitative Comparison
This section provide a quantitative comparison of the reviewed methods, focusing on their performance across different datasets. Due to the variation in the utilized datasets and the employed evaluation metrics, direct comparison between some methods are limited. Therefore, we focused to highlight the commonly reported evaluation metrics for a fair comparison, as shown in Tables 3, 4, and 5. On the CVUSA dataset, the results show CrossMLP and LGGAN achieved the best highest performance in terms of SSIM, which indicates their strong ability to generate high-quality and structurally well-defined street-view images. SSVS and PanoGAN followed closely with SSIM scores of 0.447 and 0.444, respectively, which reflects a reasonable level of structural similarity with the ground truth. In contrast, SSVG method obtained the lowest SSIM score with value of 0.3408, suggesting its poor capabilities of preserving structural elements like image texture and pattern compared to other methods. For the PSNR metric, CrossMLP reached to the highest score, followed by LGGAN with a score of 22.58. This highlights both methods’ low pixel-level reconstruction error. PGSL underperformed other methods in this metric, indicating larger difference from the ground truth. In terms of SD, LGGAN and CrossMLP methods show their high capabilities in generating clear details images with a small performance gap of 0.16. PanoGAN obtained competitive SD score, indicating it maintains comparable level of details and edge sharpness. Finally for KL score, LGGAN demonstrated its high capabilities in generating features that are close to the ground truth distribution. On the other hand, Pix2pix and X-Fork exhibited the highest KL scores, indicating that their generated images deviate from the ground truth distribution. Using the Dayton dataset, SelectionGAN and CrossMLP outperformed other methods in terms of SSIM with values of 0.593 and 0.559, respectively. These values indicate better structural similarity to the ground truth. Both methods also achieved the highest PSNR score with a small gap of 0.26, indicating low pixel-level differences with the ground truth. For SD metric, SelectionGAN achieved to the highest score, demonstrating its ability to generate sharper images with well-defined edges. Regrading the KL divergence, LGGAN outperformed other methods by achieving the lowest score, indicating its ability to generated images that are closest to the ground truth distribution and appear more realistic. SelectionGAN and CrossMLP performed well in KL metric, with differences of 0.56 and 0.15 from the best score, respectively. This small difference show that their generating images are relatively close to the ground truth distribution. In contrast, Pix2pix emerged as an outlier with the highest KL score, showing that it generates less realistic images that diverge from the original distribution. In the case of the CVACT dataset, SSVG achieved the best SSIM score with value of 0.421, revealing its strong ability to preserve the structural similarity in the generated street-view images. CrossViewDiff followed closely, highlighting comparable performance in maintaining texture and pattern quality. On the other hand, Pix2pix and X-Fork underperformed in preserving the structural elements compared to the other methods. For PSNR, SSVG and Pix2pix achieved the highest scores with a minimal gap of 0.18, implying that these two methods generate street-view images with low pixel-level errors and high visual quality. For SD, SSVG outperformed others in generating sharpest images with clear details. To ensure consistency across dataset results tables, the KL column is included for CVACT in Table 5. However, KL scores were not reported [42] and [26] from which the results were obtained, so the column is left empty.
The quantitative results of CVUSA dataset in this section, obtained from several references: Pix2pix, X-Fork, X-Seq, SelectionGAN, and PAGAN taken from [54]. SSVS and PGSL taken from [48]. CrossMLP from [38], PGSL and SSVS are from [48]. SSVG from [42], LGGAN from [47], PanoGAN from [55], and CrossViewDiff from [26]. For Dayton dataset, Pix2pix, X-Fork, X-Seq, SelectionGAN, and CrossMLP taken from [38], while LGGAN taken from [47]. For CVACT dataset, Pix2pix, X-Fork, SSVG results are taken from [42], and CrossViewDiff taken from [26]
Table 3: Quantitative comparisons on the CVUSA dataset.
SSIM $\uparrow$ PSNR $\uparrow$ SD $\uparrow$ KL $\downarrow$
Method
Pix2pix [18] 0.3847 18.0329 18.1152 14.37±1.74
X-Fork [36] 0.3698 18.6153 17.7286 14.37±1.74
X-Seq [36] 0.3737 18.9364 17.7544 10.72±1.69
SelectionGAN [46] 0.4172 19.3237 18.7789 9.18±1.63
CrossMLP [38] 0.5251 23.1532 19.58 2.69 ± 0.94
SSVS [48] 0.447 13.895 15.221 3.59 ± 0.92
PAGAN [54] 0.4315 19.4518 18.4489 7.02±1.62
SSVG [42] 0.3408 13.77 16.27 -
PGSL [57] 0.414 11.502 10.631 -
LGGAN [47] 0.5238 22.5766 19.7440 2.55 ± 0.95
PanoGAN [54] 0.4437 20.9467 19.0913 4.20±1.19
CrossViewDiff [26] 0.371 12.00 16.31 -
Table 4: Quantitative comparisons on the Dayton dataset.
SSIM $\uparrow$ PSNR $\uparrow$ SD $\uparrow$ KL $\downarrow$
Method
Pix2pix [18] 0.4180 17.6291 19.2821 38.26 ± 1.88
X-Fork [36] 0.4963 19.8928 19.4533 6.00 ± 1.28
X-Seq [36] 0.5031 20.2803 19.5258 5.93 ± 1.32
SelectionGAN [46] 0.5938 23.8874 20.0174 2.74 ± 0.86
LGGAN [47] 0.5457 22.9949 19.6145 2.18 ± 0.74
CrossMLP [38] 0.5599 23.6232 19.6688 2.33 ± 0.80
Table 5: Quantitative comparisons on the CVACT dataset.
SSIM $\uparrow$ PSNR $\uparrow$ SD $\uparrow$ KL $\downarrow$
Method
Pix2pix [18] 0.373 14.47 16.06 -
X-Fork [36] 0.370 14.17 15.89 -
SSVG [42] 0.421 14.65 16.33 -
CrossViewDiff [26] 0.412 12.411 16.29 -
7 Guidelines for Datasets Creation
This section presents guidelines for creating high- quality and diverse satellite-street-view datasets that address the challenges mentioned in Section 8.
  • Collect high quality satellite and street-view image pairs, discarding images with poor contrast, blurriness, and distortion.
  • Store metadata about the satellite and street-view pairs, including latitude and longitude, providing additional contextual information that support synthesis and geo-localization tasks.
  • Collect satellite and street-view image pairs with their corresponding auxiliary information such as segmentation maps, text prompt, satellite images with LiDAR, and 3D point cloud. These additional input will enhance the street-view generation process and maintaining geometric consistency.
  • Collect satellite and street-view image pairs from a wide geographical area, including developed and developing regions. Expanding data collection beyond the main cities helps to ensure data diversity and helps to minimize the data disparities.
  • Within city, ensure the dataset capture diverse locations, extending the data collection beyond the famous landmarks and mains roads, to capture various street patterns.
  • Ensure the collected dataset contains a diverse and balanced distribution of images across environmental settings (rural, urban and sub-urban) to minimize the data biases and improve generalization and robustness of the deep learning models.
  • Consider collecting images with time-of-day variation and under different weather conditions (e.g. rain, snow, and spring-time). Additionally, including dynamic scenes like pedestrians vehicles or traffic will help to enrich dataset diversity and ensure model robustness.
  • Develop open-access benchmark datasets is imperative for advancing satellite to street-view domain. It facilitates objective performance evaluation of deep learning models, reproducibility and fair comparison between deep learning models. In addition, publicly available datasets attract contributions from wider research community, offering interdisciplinary collaboration, resulting in the enrichment of satellite to street-view research area.
  • Create synthetic datasets that simulate extreme conditions such as occlusion, low-light environments, and noise, to effectively evaluate the robustness of developed models under these challenging real-life scenarios.
In addition to the discussed guidelines, fostering collaborations between academic institutions and organizations, such as Google Maps and Open Street Map, plays an important role in enhancing the datasets diversity and accessibility. These collaborations provide researchers with access large-scale, high-quality, and real-world datasets. Also, it helps to reduce the cost and challenges related to datasets collections.
8 Ethical and Privacy Considerations
Street-view images have becomes an important source of data in several studies and plays an important role for several applications such as urban planning and autonomous driving. However, as it uses a real-world data, it arises ethical consideration and privacy concerns related to street-view usage and synthesis. In this section we presents some of ethical and privacy concerns related to both real and synthesized street-view images. Street-view images capture real-life information that contains private properties, and potential sensitive locations. The presence of such image in datasets without owners’ consent arises ethical concerns about using these images in commercials or research domains. Also, street-view images, whether real or synthesized, can be misused in several contexts such as used for surveillance. It can be integrated with some other data sources, to construct detailed profiles about people, leading to privacy violation and misuse beyond their intended purposes. Moreover, the synthetic street-view images could be used for misinformation and fake mapping by creating non-existing locations to mislead users — for example, in real-estate and news reporting. From the privacy perspective, street-view images expose private information include individuals’ faces, private residences, license plate, and potential sensitive locations. Even though the real street-view images blur some sensitive information, some synthetics street-view images are generated with high realism, accurately generating realistic images of people or vehicles, leading to privacy violations.Furthermore, once street-view images are combined with other data such as satellite images and home addresses, they can open opportunities for criminals to plan and identify the weak security points in business and homes. To address these ethical and privacy concerns, the following strategies should be considered:
  • Advanced Anonymization: ensure an affective and advanced methods are used to obscure and blur sensitive areas and identifiable information such as not limited to, individuals’ faces, license plates, and home addresses.
  • Data Usage Consent: establishing clear consent protocol is important, to ensure that individuals are aware about how their data is being used.
  • Regularity Framework: encourage the development of guidelines and policies that govern the use and collect of street-view images . This include compliance with legal regulations such as GDPR (EU) and CCPA (California), government regulations, and regulation that criminalized the use of fake geographic data for deception.
  • Users reporting and Awareness: Enable users to report fake and specious street-view images or any geographic data to prevent fraud.
9 Real-World Applications
Street-view synthesis is useful in several tasks across different domains, including the cross-view geo-localization task, which aims to localize the street-view image by matching it against a collection of geo-referenced satellite images. It plays an important role in numerous applications like augmented reality, autonomous driving, event detection, and robotics [48],[51]. Since the retrieval component is only able to recognize and locate similar objects, integrating satellite-to-street-view synthesis plays an important role in this task. This helps the generator learn more useful information that enhances the location retrieval accuracy [48]. Also, street-view plays an important role in developing sidewalk GIS data effectively, which is considered a main component in smart city development. To utilize the sidewalk in different smart city applications, researchers require publicly accessible sidewalk GIS data that offers information about the sidewalk’s dimensions, coverage, location, and physical conditions [20], [5]. However, many cities lack such publicly available GIS data for sidewalks. The street-view images have been used for extracting the sidewalk information in the sidewalk detection task, as in [20], while both satellite and street-view images are used in the sidewalk extraction task, like [33]. Moreover, street-view images showed their importance in urban analytics and urban planning domains, as they have been used in evaluating urban infrastructure and services [66] and detecting urban changes [6].
Deploying satellite to street-view synthesis to resource-contained environments, such as mobile devices and autonomous vehicles, introduces several computational challenges due to limited battery life, computational capabilities, memory, and latency in real-time processing. To addressed these challenges several techniques need to be considered to ensure these developed models can be used effectively in resources-contained environments. The First one is employing light weight architectures, like smaller version of vision transformers (MobileVit), while maintaining high-quality images synthesis. The second one is applying model compression techniques such as quantization, pruning and knowledge distillation. The quantization technique improves the model efficiency by reducing the parameter precision, which minimized the memory usage and speeding up computation. The pruning technique eliminates less important parameters, maintaining model performance while reducing the computational costs. In addition, knowledge distillation technique enables transfer learning from well-trained, large model to a smaller one, allows the smaller model ”student” to mimic the performance of the larger one ”teacher” [11]. Moreover, hardware optimization techniques can be applied, such as running cloud-based inference for intensive processes, while running the lighter tasks locally.
10 Challenges and Limitations
@@
The following sections highlight the challenges and limitations in the Satellite to Street View generation task.
@@
Lack of multi-modality datasets: Using a single data modality depending on a single image, like a satellite image, for generating street view is more challenging than using multiple data modalities. However, some of the recent works [47, 46] rely on some auxiliary information like segmentation maps for generating street-view. Regardless, generating segmentation maps may limit the methods if there is any difficulty with its computation, affecting the quality of the synthesized street-view image. Multi-modal datasets might solve the problem and increase the quality of the generation.
Datasets Bias: satellite to street-view datasets discussed in this paper shows some biases, including limited geographical coverage and some environment setting are underrepresented. For example, CVUSA dataset focus on rural areas, where urban areas are underrepresented. The CVACT dataset covers wider range of urban and sub-urban areas compared to CVUSA [53], where Dayton dataset [50] cover more urban areas.These biases in the datasets within this domain results in a challenge for training of deep learning models. Once these models trained on dataset biased to specific region, the learned pattern will be limited to those areas, leading to challenges in capturing the minority environment settings.
@@
Lack of novel techniques: The methods used in the literature need to be updated; currently, generative adversarial network and their variants are the most popular choices for a street-view generation. Moreover, various CNNs are employed for feature extraction, while Siamese and Triplet Networks have been used for street-view geo-localization tasks.
@@
One of the key challenges for generative models is maintaining robustness against noisy inputs, which can significantly degrade the quality of extracted features. In many real-world domains, input data are often corrupted due to environmental factors. This issue is particularly prominent in satellite imagery, where noise can result from factors such as cloud cover, sensor defects, and adverse weather conditions. Such noise can have a substantial negative impact on performance in tasks like image synthesis. Therefore, enhancing model robustness to noisy inputs is critical for ensuring the reliability and quality of the generated outputs, as well as improving generalization across diverse real-world scenarios.
Several strategies can be employed to address the issue of noisy inputs. One approach involves applying deep learning-based pre-processing techniques to denoise inputs before they are fed into generative models. For example, Denoising Autoencoders (DAEs) [49] are trained to reconstruct clean images from noisy versions. Additionally, incorporating auxiliary information—such as geo-spatial metadata or semantic maps—can provide valuable complementary signals when satellite imagery is partially obscured. For instance, when cloud cover limits visual clarity, auxiliary data can help infer the underlying scene. Furthermore, data augmentation that includes noisy samples during training can improve model robustness and generalization, effectively preparing the model to handle noisy data encountered in real-world applications.
11 Future Research Directions
Despite of progress made in satellite to street-view synthesis domain, there are still some challenges that need to be addressed. To further advance this field, these challenges need to be addressed and future research should focus on the following:
  • Researchers must consider and focus on recent state-of-the-art techniques such as transformers and stable diffusion for satellite to street-view synthesizing.
  • Integrate multi-modal data such as depth maps, text description, and LiDAR to provide richer spatial and contextual information, resulting in improving model accuracy and enhances its generalization capabilities.
  • In addressing the lack of multi-modalities, integrating Multi-interactive Feature Learning mechanisms could improve the model’s ability to leverage complementary information across data sources (e.g., satellite imagery, semantic maps, and LiDAR). Moreover, the creation of a Full-time Multi-modality Benchmark—one that includes synchronized multi-modal data under diverse real-world conditions—would provide a standardized foundation for evaluating image fusion and segmentation methods.
  • For the challenge of low-resolution layout images, incorporating super-resolution techniques presents a promising avenue. These methods can be applied either as pre-processing steps or embedded within generative frameworks to enhance spatial detail, which is crucial for downstream tasks such as fine-grained localization and image synthesis. Exploring cross-domain super-resolution approaches, particularly those adapted to the unique characteristics of satellite and layout imagery, may significantly improve model performance in this context.
  • Create high-quality, large-scale, divers, and open-access datasets, as recommended Section 5 (Guidelines of Datasets Creation), to enrich research in this area.
  • Employ cross-dataset training, to expose model to a divers environmental settings, which improves the model’s generalization.
  • Transformers and diffusion model have shown promising results in satellite to street-view task. In the future, leveraging these models or their variants more extensively in this domain offer valuable opportunities. Also, integrating Large Language model (LLMs) with the diffusion models could further enhance the generation process and helps to get enrich information and generate a descriptive prompt for satellite images to guide the diffusion model. This integration can help to provide contextual understanding and leading to generating more realistic and accurate street-view images. The diffusion models are capable of generating street-view images with high-quality and fine-grained details. However, they sometimes generate a bit generic images and may lack full geometric consistency with the ground truth. Adopting a hybrid approach that combines diffusion model and GAN could help to improve the overall framework by leveraging the strengths of both models. This incorporation enables to learn more robust features, resulting in improved street-view generation.
12 Conclusion
This paper has surveyed the satellite to street-view synthesis domain from different aspects, including the recently available methods, common datasets (benchmark datasets), common evaluation metrics, challenges and limitations, by examining 23 recently published papers. It is clear from the reviewed papers that the Generative Neural Network (GAN) and its varieties are very immersed and widely used in this domain, despite other novel methods being available, diffusion and transformer models have received little attention in this domain. Despite the success of the reviewed methods, they still struggle to generate small and detailed objects, which confirms that novel methods need to be used to overcome these limitations and generate more realistic and detailed street-view images. In addition, the research highlights that insufficient diverse publicly available datasets, and most of the surveyed papers rely on limited datasets. This indicates the need for more high-quality, publicly available datasets to enrich the research in this domain. Moreover, methods in the literature have been evaluated on image quality-based metrics, indicating the importance of more specific evaluation metrics to evaluate the generated images. Overall, the disclosed research challenges highlight that the satellite to street-view synthesis is still an active area of research. Considering the interdisciplinary nature of this field, effectively addressing these challenges will require cross-disciplinary collaborations between geospatial science, urban planning, and machine learning to foster the development of advanced methods and utilize diverse datasets. The developments in satellite to street-view synthesis domain offer considerable policy implications in several domains, including environmental management, urban planning, disaster response and smart cities. By leveraging up-to-date and high-resolution data, policymakers can gain valuable insights and improve the decision-making processes in areas like infrastructure evaluation and climate change adaptation. It is essential for academic institutions, governments, and industries to collaborate in advancing and deploying these models effectively in real-world applications. We suggest to integrating this domain into education curricula to prepare future researchers and train them in geospatial AI, for advancing this domain. This integration will equip researchers with the required skills to address challenges, ultimately enhancing applications in several domains such as urban planning and autonomous driving. We also encourage researchers to contribute in addressing the outlined challenges and to foster research collaborations to drive progress and innovations in this research area.
Acknowledgement: A preprint of this paper is available on arXiv at https://arxiv.org/abs/2405.08961. The author/s would like to acknowledge the support received from Saudi Data and AI Authority (SDAIA) and King Fahd University of Petroleum and Minerals (KFUPM) under SDAIA-KFUPM Joint Research Center for Artificial Intelligence Grant No. JRC-AI-RFP-11.